OpenStudy (anonymous):

2x+8=14 what is the value of x in this question

OpenStudy (anonymous):

Solve it and see? What would you do first to get the x term by itself?

OpenStudy (anonymous):

idk how to do it

OpenStudy (anonymous):

How do you solve something like 4+x = 5?

OpenStudy (anonymous):

put the 4 on both sides

OpenStudy (anonymous):

I think you mean subtract 4 from both sides?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

And what do you get when you do that?

OpenStudy (anonymous):

one gonna be zero than the other side will be 1

OpenStudy (anonymous):

Not quite \(4+x = 5\) -4 -4 \(0 + x = 1\) \[\implies x = 1\]

OpenStudy (anonymous):

So now how would you solve this one (using the same method) k + 8 = 14

OpenStudy (anonymous):

add eight on both sides

OpenStudy (anonymous):

subtract 8. You want the 8 on the left to become a 0

OpenStudy (anonymous):

not add.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

So what do you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Did you see what I did before?

OpenStudy (anonymous):

Cause you need to be able to do it yourself or things are only going to get harder from here.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

Are you?

OpenStudy (anonymous):

yea